OpenStudy (nbarrera):

Help Me+Medal! Arrange the following fractions in order from least to greatest.

OpenStudy (nbarrera):

\[\frac{1 }{ 3 },\frac{ 2 }{ 5 },\frac{ 3 }{ 8 }\]

OpenStudy (texaschic101):

with problems like these, you can either turn them all into fractions with the same denominators and then compare them....or, turn them all into decimals and compare them.

OpenStudy (texaschic101):

from the looks of the fractions, decimals will be easier.. 1/3 = 0.3333 2/5 = 0.4 3/8 = 0.375 so what do you think the order is now ? from least to greatest

OpenStudy (nbarrera):

So 0.3333, 0.375, 0.4

OpenStudy (texaschic101):

yes....1/3, 3/8, 2/5